# HMARL-CBF – Hierarchical Multi-Agent Reinforcement Learning with Control Barrier Functions for Safety-Critical Autonomous Systems

### v1

@@ 6.4 Results and Baseline Comparison
[…]
@@ 6.4.1 Evaluation Setup and Results
  Merging Success $\uparrow$ $0.43_{\pm 0.02}$ $0.32_{\pm 0.04}$ $0.43_{\pm 0.09}$ $0.48_{\pm 0.01}$     $0.99_{\pm 0.01}$ $0.99_{\pm 0.00}$ $0.99_{\pm 0.01}$ $0.99_{\pm 0.01}$
@@
Energy $\downarrow$ $5.90_{\pm 0.62}$ $6.64_{\pm 0.56}$ $6.52_{\pm 0.27}$ $5.68_{\pm 0.39}$     $5.56_{\pm 0.04}$ $3.62_{\pm 0.68}$ $5.78_{\pm 0.24}$ $5.11_{\pm 0.20}$
Roundabout Success $\uparrow$ $0.85_{\pm 0.05}$ $0.82_{\pm 0.04}$ $0.85_{\pm 0.04}$ $0.80_{\pm 0.03}$     $0.99_{\pm 0.01}$ $0.99_{\pm 0.04}$ $0.98_{\pm 0.03}$ $0.98_{\pm 0.02}$
@@
Energy $\downarrow$ $11.74_{\pm 0.52}$ $12.69_{\pm 0.18}$ $12.07_{\pm 0.67}$ $12.31_{\pm 0.26}$     $9.89_{\pm 0.23}$ $7.75_{\pm 0.39}$ $10.15_{\pm 0.82}$ $11.41_{\pm 0.69}$
Bottleneck Success $\uparrow$ $0.09_{\pm 0.03}$ $0.34_{\pm 0.15}$ $0.48_{\pm 0.16}$ $0.75_{\pm 0.07}$     $0.99_{\pm 0.00}$ $0.99_{\pm 0.02}$ $0.99_{\pm 0.01}$ $0.99_{\pm 0.01}$
@@
Tollgate Success $\uparrow$ $0.08_{\pm 0.37}$ $0.05_{\pm 0.02}$ $0.11_{\pm 0.04}$ $0.00_{\pm 0.00}$     $0.99_{\pm 0.02}$ $0.98_{\pm 0.02}$ $0.98_{\pm 0.01}$ $0.99_{\pm 0.01}$
Time $\downarrow$ $2809.04_{\pm 364.04}$ $3211.06_{\pm 511.27}$ $1884.36_{\pm 319.82}$     $369.32_{\pm 5.82}$ $427.02_{\pm 32.43}$ $377.42_{\pm 9.99}$ $307.82_{\pm 8.09}$
Energy $\downarrow$ $74.64_{\pm 10.47}$ $114.25_{\pm 27.02}$ $53.63_{\pm 7.72}$     $8.40_{\pm 0.05}$ $5.17_{\pm 0.44}$ $8.93_{\pm 0.11}$ $8.99_{\pm 0.01}$
@@ 6.5 Ablation Results
Merging Success $\uparrow$ 0 0 0 $0.94_{\pm 0.09}$ $0.91_{\pm 0.04}$ 0 $0.52_{\pm 0.06}$ $0.99_{\pm 0.01}$
Energy $\downarrow$ - - - $14.17_{\pm 0.49}$ $15.30_{\pm 0.31}$ $22.30_{\pm 0.12}$ $14.10_{\pm 0.23}$
Time $\downarrow$ - - - $211.70_{\pm 23.80}$ $183.91_{\pm 1.49}$ $164.52_{\pm 2.64}$
Intersection Success $\uparrow$ 0 0 0 $0.86_{\pm 0.04}$ $0.80_{\pm 0.04}$ $0.14_{\pm 0.02}$ 0 $0.97_{\pm 0.04}$
Energy $\downarrow$ - - - $5.75_{\pm 0.08}$ $5.00_{\pm 0.11}$ $50.00_{\pm 0.71}$ $5.56_{\pm 0.04}$
Time $\downarrow$ - - - $574.30_{\pm 60.60}$ $320.50_{\pm 68.17}$ $921.42_{\pm 40.71}$ $317.41_{\pm 33.92}$
Roundabout Success $\uparrow$ 0 0 0 $0.68_{\pm 0.08}$ $0.60_{\pm 0.08}$ 0 0 $0.99_{\pm 0.01}$
Energy $\downarrow$ - - - $13.23_{\pm 1.08}$ $8.50_{\pm 0.74}$ $9.89_{\pm 0.23}$
Time $\downarrow$ - - - $907.35_{\pm 50.98}$ $600.03_{\pm 34.67}$ $403.18_{\pm 21.66}$
Bottleneck Success $\uparrow$ 0 0 0 $0.65_{\pm 0.03}$ $0.55_{\pm 0.09}$ 0 0 $0.99_{\pm 0.00}$
Energy $\downarrow$ - - - $10.03_{\pm 0.10}$ $12.03_{\pm 0.04}$ $7.37_{\pm 0.00}$
Time $\downarrow$ - - - $923.07_{\pm 61.73}$ $965.45_{\pm 24.18}$ $241.16_{\pm 8.33}$
Tollgate Success $\uparrow$ 0 0 0 $0.91_{\pm 0.03}$ $0.92_{\pm 0.03}$ 0 0 $0.99_{\pm 0.02}$
Energy $\downarrow$ - - - $8.68_{\pm 0.03}$ $9.35_{\pm 0.02}$ $8.40_{\pm 0.05}$
Time $\downarrow$ - - - $713.00_{\pm 24.15}$ $423.40_{\pm 15.60}$ $369.32_{\pm 5.82}$
@@ 6.5.1 Generalization Results
  Lane Width (m) Success Rate Merging Intersection Roundabout Bottleneck Tollgate   3.25 $0.99_{\pm 0.01}$ $0.90_{\pm 0.03}$ $0.98_{\pm 0.02}$ $0.97_{\pm 0.02}$ $0.99_{\pm 0.001}$ 3.5 (default) $0.99_{\pm 0.01}$ $0.97_{\pm 0.04}$ $0.99_{\pm 0.01}$ $0.99_{\pm 0.003}$ $0.99_{\pm 0.02}$ 4 $0.99_{\pm 0.01}$ $0.97_{\pm 0.015}$ $0.99_{\pm 0.01}$ $0.99_{\pm 0.001}$ $0.99_{\pm 0.02}$
@@ 6.6.1 Implementation Details: METADRIVE Simulator
$r^{i}_{\mathrm{L}}=-\,c_{1}||\bm{a}^{i}||^{2}-c_{2}\!\left(\frac{v^{i}-v_{\mathrm{des}}}{v_{\mathrm{des}}}\right)^{\!2}-c_{3}\!\left(\frac{\psi^{i}-\psi_{\mathrm{des}}}{\pi}\right)^{\!2}-c_{4}||d^{i}||^{2}$ (39)
@@ 6.6.2 Implementation Details: Lidar Suite Simulator
Target. Agents dynamics in this environment are governed by double integrator dynamics with continuous-time states $\bm{x}_{i}=[p_{x}^{i},p_{y}^{i},v_{x}^{i},v_{y}^{i}]^{\top}\in\mathbb{R}^{4}$, where $\bm{p}_{i}=[p_{x}^{i},p_{y}^{i}]^{\top}\in\mathbb{R}^{2}$ denotes position and $\bm{v}_{i}=[v_{x}^{i},v_{y}^{i}]^{\top}\in\mathbb{R}^{2}$ denotes velocity; control inputs $\bm{u}_{i}=[a_{x}^{i},a_{y}^{i}]^{\top}\in\mathbb{R}^{2}$ represent accelerations along the Cartesian axes, and the agent dynamics follow $\dot{\bm{x}}_{i}=[v_{x}^{i},v_{y}^{i},a_{x}^{i},a_{y}^{i}]^{\top}$, with velocity and control constraints bounded in $[-10,10]$ and $[-1,1]$ respectively.
Spread. The underlying dynamics, state representation, control parametrization, and bounds are identical to those in the Target environment,i.e., agents evolve under double integrator dynamics with four-dimensional state vectors comprising positions and velocities, and two-dimensional control inputs that modulate accelerations; the bounded constraints on velocities and controls likewise remain $[-10,10]$ and $[-1,1]$, respectively.

### v2

@@ 6.4 Results and Baseline Comparison
We trained with 15 vehicles for merging, 30 vehicles for intersection and roundabout, 25 vehicles for bottleneck and 40 vehicles for tollgate respectively. As mentioned previously, the extrinsic reward corresponding to (14) was set to the environment reward, besides the regularization terms. The intrinsic reward is described in subsection 6.6. We present the results of training for the Merging, Bottleneck and Tollgate environments in Figures 4 and 5 respectively. The key observation is that our algorithm achieves almost 100 $\%$ success rate with an approximate model of the vehicle (see appendix 6.6) from the beginning of training, while also optimizing the joint policy evident from the average episode length. Although several baselines achieve better average episode length in the tollgate environment, the success rate is very low, implying most agents crashes out resulting in the remaining agents to finish faster, returning a lower average episode length. With the hierarchical decomposition, our algorithms converge faster ($\leq$ 300k) compared to the baselines which we run for 1M iterations, thus confirming the claim about expediting convergence.
@@ 6.4.1 Evaluation Setup and Results
  Merging Success $\uparrow$ $0.43_{\pm 0.02}$ $0.32_{\pm 0.04}$ $0.43_{\pm 0.09}$ $0.48_{\pm 0.01}$     $9_{\pm 0.01} 0.99±0$$9_{\pm 0.00}$ $9_{\pm 0.01} 0.99±0$$9_{\pm 0.01}$
@@
Energy $\downarrow$ $5.90_{\pm 0.62}$ $6.64_{\pm 0.56}$ $6.52_{\pm 0.27}$ $68_{\pm 0.39}     5.5$$56_{\pm 0.04}$ $3.62_{\pm 0.68}$ $5.78_{\pm 0.24}$ $5.11_{\pm 0.20}$
Roundabout Success $\uparrow$ $0.85_{\pm 0.05}$ $0.82_{\pm 0.04}$ $0.85_{\pm 0.04}$ $0.80_{\pm 0.03}$     $9_{\pm 0.01} 0.99±0$$9_{\pm 0.04}$ $0.98_{\pm 0.03}$ $0.98_{\pm 0.02}$
@@
Energy $\downarrow$ $11.74_{\pm 0.52}$ $12.69_{\pm 0.18}$ $12.07_{\pm 0.67}$ $12.31_{\pm 0.26}$     $9.89_{\pm 0.23}$ $7.75_{\pm 0.39}$ $0.15_{\pm 0.82} 11$$1.41_{\pm 0.69}$
Bottleneck Success $\uparrow$ $0.09_{\pm 0.03}$ $0.34_{\pm 0.15}$ $0.48_{\pm 0.16}$ $0.75_{\pm 0.07}$     $9_{\pm 0.00} 0.99±0$$9_{\pm 0.02}$ $9_{\pm 0.01} 0.99±0$$9_{\pm 0.01}$
@@
Tollgate Success $\uparrow$ $08_{\pm 0.37} 0.0$$05_{\pm 0.02}$ $11_{\pm 0.04} 0.0$$00_{\pm 0.00}$     $0.99_{\pm 0.02}$ $0.98_{\pm 0.02}$ $8_{\pm 0.01} 0.99±0$$9_{\pm 0.01}$
Time $\downarrow$ $2809.04_{\pm 364.04}$ $3211.06_{\pm 511.27}$ $1884.36_{\pm 319.82}$ -     $369.32_{\pm 5.82}$ $427.02_{\pm 32.43}$ $377.42_{\pm 9.99}$ $307.82_{\pm 8.09}$
Energy $\downarrow$ $74.64_{\pm 10.47}$ $114.25_{\pm 27.02}$ $53.63_{\pm 7.72}$ -     $8.40_{\pm 0.05}$ $5.17_{\pm 0.44}$ $8.93_{\pm 0.11}$ $8.99_{\pm 0.01}$
@@ 6.5 Ablation Results
[…]
Energy $\downarrow$ - - - $14.17_{\pm 0.49}$ $15.30_{\pm 0.31}$ - $22.30_{\pm 0.12}$ $14.10_{\pm 0.23}$
Time $\downarrow$ - - - $211.70_{\pm 23.80}$ $183.91_{\pm 1.49}$ - - $164.52_{\pm 2.64}$
[…]
Energy $\downarrow$ - - - $5.75_{\pm 0.08}$ $5.00_{\pm 0.11}$ $0.00_{\pm 0.71} - 5.5$$56_{\pm 0.04}$
Time $\downarrow$ - - - $574.30_{\pm 60.60}$ $320.50_{\pm 68.17}$ $921.42_{\pm 40.71}$ - $317.41_{\pm 33.92}$
Roundabout Success $\uparrow$ 0 0 0 $0.68_{\pm 0.08}$ $60_{\pm 0.08} 0 0$ $0.99_{\pm 0.01}$
Energy $\downarrow$ - - - $13.23_{\pm 1.08}$ $8.50_{\pm 0.74}$ - - $9.89_{\pm 0.23}$
Time $\downarrow$ - - - $907.35_{\pm 50.98}$ $600.03_{\pm 34.67}$ - - $403.18_{\pm 21.66}$
Bottleneck Success $\uparrow$ 0 0 0 $0.65_{\pm 0.03}$ $55_{\pm 0.09} 0 0$ $0.99_{\pm 0.00}$
Energy $\downarrow$ - - - $10.03_{\pm 0.10}$ $12.03_{\pm 0.04}$ - - $7.37_{\pm 0.00}$
Time $\downarrow$ - - - $923.07_{\pm 61.73}$ $965.45_{\pm 24.18}$ - - $241.16_{\pm 8.33}$
Tollgate Success $\uparrow$ 0 0 0 $0.91_{\pm 0.03}$ $92_{\pm 0.03} 0 0$ $0.99_{\pm 0.02}$
Energy $\downarrow$ - - - $8.68_{\pm 0.03}$ $9.35_{\pm 0.02}$ - - $8.40_{\pm 0.05}$
Time $\downarrow$ - - - $713.00_{\pm 24.15}$ $423.40_{\pm 15.60}$ - - $369.32_{\pm 5.82}$
@@ 6.5.1 Generalization Results
  Lane Width (m) Success Rate Merging Intersection Roundabout Bottleneck Tollgate   3.25 $0.99_{\pm 0.01}$ $0.90_{\pm 0.03}$ $0.98_{\pm 0.02}$ $7_{\pm 0.02} 0.99±0$$9_{\pm 0.001}$ 3.5 (default) $0.99_{\pm 0.01}$ $7_{\pm 0.04} 0.99±0$$9_{\pm 0.01}$ $9_{\pm 0.003} 0.99±0$$9_{\pm 0.02}$ 4 $0.99_{\pm 0.01}$ $7_{\pm 0.015} 0.99±0$$9_{\pm 0.01}$ $9_{\pm 0.001} 0.99±0$$9_{\pm 0.02}$
@@ 6.6.1 Implementation Details: METADRIVE Simulator
[…]
@@ 6.6.2 Implementation Details: Lidar Suite Simulator
Target. Agents dynamics in this environment are governed by double integrator dynamics with continuous-time states $\bm{x}_{i}=[p_{x}^{i},p_{y}^{i},v_{x}^{i},v_{y}^{i}]^{\top}\in\mathbb{R}^{4}$, where $\bm{p}_{i}=[p_{x}^{i},p_{y}^{i}]^{\top}\in\mathbb{R}^{2}$ denotes position and $\bm{v}_{i}=[v_{x}^{i},v_{y}^{i}]^{\top}\in\mathbb{R}^{2}$ denotes velocity; control inputs $\bm{u}_{i}=[a_{x}^{i},a_{y}^{i}]^{\top}\in\mathbb{R}^{2}$ represent accelerations along the Cartesian axes, and the agent dynamics follow $\dot{\bm{x}}_{i}=[v_{x}^{i},v_{y}^{i},a_{x}^{i},a_{y}^{i}]^{\top}$, with velocity and control constraints bounded in [10,10][-10,10] and [1,1][-1,1] respectively.
Spread. The underlying dynamics, state representation, control parametrization, and bounds are identical to those in the Target environment,i.e., agents evolve under double integrator dynamics with four-dimensional state vectors comprising positions and velocities, and two-dimensional control inputs that modulate accelerations; the bounded constraints on velocities and controls likewise remain [10,10][-10,10] and [1,1][-1,1], respectively.